# Asynchronous expressed-private $q$-voter model on networks: self-anticonformity and preference falsification

### v2

@@ Abstract
People may express preferences that differ from their privately held views, often under social pressure, and may fail to act on their stated intentions. Such inconsistencies are referred to as preference falsification and the intention-behavior gap, respectively. Both hamper collective decision-making and adaptation, complicating policy formulation and implementation. To simulate these phenomena, dual-layer opinion agent-based models are used, in which each agent holds both a private and an expressed (public) opinion. Within the $q$-voter framework, two expressed-private opinion (EPO) models have been introduced in which private and expressed opinions are updated synchronously; two variants differ only by the presence of self-anticonformity, a mechanism in which an agent may set its private opinion opposite to its current expressed opinion, breaking internal harmony and creating a state akin to cognitive dissonance. Here, we extend these models by introducing an asynchronous update: in each elementary step, an agent updates its private opinion with probability $\alpha$ or its expressed opinion with complementary probability; hence the name $\alpha$-EPO models. Using Monte Carlo simulations on both artificial and real organizational social networks, along with mean-field and pair approximation analyses, we show that self-anticonformity makes collective outcomes robust to behavioral volatility tuned by $\alpha$, enhances collective agreement, and suppresses hysteresis. In contrast, without self-anticonformity, $\alpha$ affects the nature of the transition between agreement and disagreement: higher values of $\alpha$ suppress hysteresis and enhance overall agreement.
preprint: AIP/123-QED
@@ I Introduction
Examples of discrepancies between what people think and what they openly say, or between what they say and what they do, are common across all spheres of life. For instance, a large body of literature shows a persistent gap between consumers’ stated concern for the environment, on the one hand, and their actual behaviors aimed at environmental sustainability, on the other hand, as reviewed in ElHaffar et al. (2020). Moreover, it is well known that people do not always express their true preferences, a phenomenon known as preference falsification, because they may fear social disapproval, sanctions, or negative consequences arising from openly stating views that deviate from perceived social norms Kuran (1987). A natural modeling approach to such situations is to distinguish between an agent’s private and expressed states. This idea underlies the class of expressed-private opinion (EPO) models Dong et al. (2024); for a short and recent review, see Kamińska and Sznajd-Weron (2025).
@@
Here, we focus on a specific aspect of these models: the effect of unequal update rates for private and expressed opinions. Some EPO-type models assume that private opinions do not change over time Masuda and Redner (2011); Gaisbauer et al. (2020); Manfredi et al. (2020); Llabrés et al. (2023). By contrast, others assume that behavior and attitudes may change with equal rates Banisch and Olbrich (2019); Ye et al. (2019); León-Medina et al. (2020); Jacob and Banisch (2023); Dong et al. (2024); Peng et al. (2025). Finally, there are models in which the update rates for expressed and private opinions differ. An interesting example is the model proposed by Piras et al. (2022), which assumes that agents are generally more likely to change their behavior than their attitudes. When agents experience strong cognitive dissonance Festinger (1957), however, whether they change behavior or opinion depends on how often they have altered their opinions in the past. Another example, much closer to what we propose in this paper, is the concealed voter model Gastner et al. (2018), in which agents change their expressed opinion by copying the opinion of a randomly selected neighbor (with probability $c$) or externalizing their internal opinion (with probability $e$), while the private opinion is updated by internalizing the expressed one (with probability $i$).
@@
Asynchronous updating has not yet been studied in $q$-voter based EPO models. The two EPO variants introduced within this framework Jędrzejewski et al. (2018); Kamińska and Sznajd-Weron (2025) assume that private and expressed opinions are updated synchronously; in this setting, it has been shown that the stationary average private and expressed opinions (i.e., the collective-level aggregates) are independent of whether agents update their private opinion before their expressed opinion, or vice versa Jędrzejewski et al. (2018); Kamińska and Sznajd-Weron (2025). It remains unclear, however, whether introducing rate asymmetry between private and expressed opinions alters the collective dynamics. We therefore extend these models by introducing an asynchronous update scheme in which, at each elementary step, only one layer is updated: private opinions are updated with probability $\alpha$, and expressed opinions with the complementary probability $1-\alpha$. We refer to these extensions as $\alpha$-EPO models. We study both variants considered in  Jędrzejewski et al. (2018); Kamińska and Sznajd-Weron (2025) and analyze the resulting dynamics on synthetic and empirical networks using mean-field and pair approximation, as well as Monte Carlo simulations.
@@
Studying both variants in parallel is particularly interesting because, as recently suggested, the older version of the model can be interpreted as incorporating a form of self-anticonformity Jędrzejewski et al. (2018), whereas the newer version does not include this mechanism Kamińska and Sznajd-Weron (2025). Furthermore, Jędrzejewski et al. (2018) and Kamińska and Sznajd-Weron (2025) considered their models only on a complete graph, corresponding to a fully connected society in which everyone knows everyone. For large social systems, this assumption is clearly unrealistic. Here, we move beyond the complete graph by studying $\alpha$-EPO $q$-voter models on networks using Monte Carlo (MC) simulations and pair approximation (PA) analysis.
@@ II The model
We consider a system of $N$ agents, indexed by $i=1,\ldots,N$, each described by two binary variables: $S_{i}(t)\in\{-1,+1\}$, the expressed opinion of agent $i$ at time $t$ (visible to other agents), and $\sigma_{i}(t)\in\{-1,+1\}$, the private opinion of agent $i$ at time $t$ (not visible to others). The binary variables represent a choice between two alternatives, such as yes or no; for or against; adopting or not adopting, etc. Agents occupy the nodes of an undirected, unweighted network, whose topology specifies the interaction patterns within the system. For each agent $i$, let $\mathcal{N}(i)$ denote the set of its neighbors and let $k_{i}=|\mathcal{N}(i)|$ be the degree of node $i$; thus, agent $i$ can interact only with agents in $\mathcal{N}(i)$. In conformity events, the influence group (the $q$-panel) is formed by selecting uniformly at random, without replacement, $q$ neighbors of agent $i$, i.e.,a subset $\mathcal{Q}\subset\mathcal{N}(i)\;\wedge\;|\mathcal{Q}|=q$. In all analytical calculations and algorithmic description of the model we assume that every node in the network has at least $q$ neighbors.
@@
Agents can update their opinions at both levels (expressed and private) due either to social influence from their neighbors or to the need to reduce internal dissonance, which arises when private and expressed opinions are in conflict. The updating rules for public and private opinions differ; however, at both levels we consider two mechanisms: independence and conformity. Independence at the expressed level means that an agent simply verbalizes its private opinion, i.e., its expressed opinion is set equal to the private one, thereby reducing the agent’s internal dissonance. In contrast, independence at the private level corresponds to stochastic noise, as in many other versions of the nonlinear voter model Carro et al. (2016); Khalil et al. (2018); Khalil and Toral (2019); Peralta et al. (2018); Vieira et al. (2020). Specifically, we use the formulation proposed in Nyczka et al. (2012): under independence, an agent flips its private opinion to the opposite one with probability $1/2$.
@@
Conformity at both levels is driven by social pressure exerted by other agents. In our models, the source of this social influence is a group of $q$ randomly selected neighbors, called the $q$-panel. At the expressed level, conformity comprises two scenarios, depending on the agent’s initial state. If an agent is initially in harmony (i.e., its private and expressed opinions coincide), it is less susceptible to influence. Therefore, it adjusts its expressed opinion to match the group only if the $q$-panel is unanimous, that is, if all $q$ selected neighbors share the same expressed opinion. This type of behavior is called compliance Nail et al. (2013). By contrast, an agent who is in dissonance and thus experiences internal tension is more willing to change. In this case, it is enough that at least one member of the $q$-panel publicly expresses an opinion consistent with the agent’s private beliefs to encourage the agent to speak up, align its expressed opinion with its private one, and thereby relieve dissonance. This type of response to social influence is referred to as disinhibitory contagion Nail et al. (2013), a phenomenon in which breaking the norm by even a single individual incites others to act according to their true intentions.
@@
Conformity at the private level also requires a source of influence. Here, a key difference emerges between the models proposed by Jędrzejewski et al. (2018) and Kamińska and Sznajd-Weron (2025), and this distinction carries over to the two variants considered in this paper. In the first case, following Jędrzejewski et al. (2018), an agent conforms to a unanimous $q$-panel regardless of its expressed opinion. Consequently, the agent may change its private opinion to match the group while still expressing the opposite view. We refer to this as self-anticonformity, since the agent, at the expressed level, disagrees not only with the group but also with its own private opinion. In the second variant, following Kamińska and Sznajd-Weron (2025), this scenario is prohibited. In other words, an agent conforms privately only if the $q$-panel is unanimous and its opinion matches the agent’s expressed one. This also implies that conformity on the private level cannot lead to dissonance.
@@
In this paper, we introduce an extension that applies to both variants of the model described above, allowing for a more flexible relationship between updates at the expressed and private levels. The proposed extension captures the fact that a change in private opinion is not necessarily followed by a change in behavior, and vice versa: behavior may change without altering beliefs Kroesen and Chorus (2018). Both Jędrzejewski et al. (2018); Kamińska and Sznajd-Weron (2025) assumed that during a single update, both expressed and private opinions are updated. They considered two update schemes: “act then think” (AT), in which the expressed opinion is updated first and the private one second, and “think then act” (TA), in which the private opinion is updated first and the expressed one second. Here, we assume that with probability $\alpha$ an agent updates its private opinion, and with complementary probability $1-\alpha$ it updates its expressed opinion. The algorithm for an elementary update of the $\alpha$-EPO model, taking time $\Delta t$, is described by Algorithm 1 and illustrated in Figs. 1 and 2.
@@
[…]
@@
[…]
@@
Let us, following Jędrzejewski et al. (2018); Kamińska and Sznajd-Weron (2025), use $\uparrow$ to denote the opinion +1+1 and $\downarrow$ to denote the opinion 1-1. To describe the state of a single agent at time $t$, we use the ordered pair $(S(t),\sigma(t))$, where the first element is the expressed opinion, followed by the private one. Accordingly, the set of possible agent states is $X=(S,\sigma)\in\{\uparrow\uparrow,\uparrow\downarrow,\downarrow\uparrow,\downarrow\downarrow\}$. The macroscopic state of the system at time $t$ can be described by the fractions of agents in each state, denoted by $c_{X}(t)$, defined as
@@
To enable comparison with Jędrzejewski et al. (2018) and Kamińska and Sznajd-Weron (2025) instead of analyzing the concentrations of agents in all possible states separately, we consider the following quantities:
@@ III.1 Monte Carlo simulations
We use Monte Carlo simulations both to validate our analytical calculations for complete and random graphs, as well as to extend the analysis to other topologies, including the empirical networks. To obtain smooth trajectories comparable with the mean-field and pair approximations (which effectively correspond to the thermodynamic limit), we simulate relatively large systems with $N=10^{4}$ and $N=10^{5}$ nodes. We also conduct MC simulations on empirical networks from organizations Fire and Puzis (2016). In the latter case, some nodes may have degree lower than $q$. In such cases we assume that the $q$-panel consists of all neighbors, i.e., for each agent $i$ we take $q_{i}=\min(k_{i},q)$.
@@ III.2.1 Model with self-anticonformity
We can proceed analogously to Jędrzejewski et al. (2018), that is we add (14) and (15), then (14) and (16). Then, using Eqs. (3) and (4) we finally obtain
@@
Hence for $\alpha\neq 0$ and $\alpha\neq 1$ we obtain exactly the same system of equations as Jędrzejewski et al. (2018), namely
@@ III.3 Pair approximation
Pair approximation is an analytical approach that relaxes the assumption of a completely homogeneous agent system by accounting for dynamical correlations between pairs of agents in different states. Within this framework, the evolution equations for the state concentrations in the model with self-anticonformity are given by the following system (for the model without self-anticonformity, the framed terms are omitted):
@@
As in other homogeneous pair-approximation treatments of $q$-voter models Jędrzejewski and Sznajd-Weron (2022); Lipiecki and Sznajd-Weron (2025b), correlations between node degree and state do not emerge unless they are present in the initial condition. Hence, we can set $\langle k\rangle_{X}=\langle k\rangle$ for all $X$ and extract the opinion concentration from Eq. (42) according to:
@@ IV.1 Mean-field results
As shown in Fig. 3, both models exhibit a phase transition between an ordered phase (agreement), in which the majority of agents share the same opinion ($c_{S}\neq 0.5$, $c_{\sigma}\neq 0.5$), and a disordered phase (disagreement), in which opinions are evenly split ($c_{S}=0.5$, $c_{\sigma}=0.5$). The location of the critical point $p^{*}$ and the type of the transition differ between the two $\alpha$-EPO variants. Nevertheless, within each variant, $c_{S}(p)$ and $c_{\sigma}(p)$ undergo the transition at the same critical point, and the transition is of the same type for both quantities. In the ordered phase ($p<p^{*}$), the majority opinion is more pronounced at the expressed level than at the private level. This result is consistent with pluralistic ignorance, i.e., a mismatch in which privately held views are more diverse than publicly expressed ones because individuals misperceive the prevailing norm and adjust their public behavior accordingly Miller (2023).
@@
The role of $\alpha$ differs qualitatively between the two variants of the $\alpha$-EPO model. In the version with self-anticonformity, the stationary phase diagrams $c_{S}(p)$ and $c_{\sigma}(p)$ are independent of $\alpha$ and therefore coincide with those reported in Jędrzejewski et al. (2018) (top row in Fig. 3). This is consistent with Eqs. (20)–(21), which, for $0<\alpha<1$, do not contain $\alpha$. In this variant, $\alpha$ affects only the stationary dissonance $d$, which increases with $\alpha$. This agrees with Jędrzejewski et al. (2018), where dissonance was also the only quantity that distinguished the two update schemes: act then think (AT), in which the expressed opinion was updated first and the private opinion second, and think then act (TA), in which the private opinion was updated first and the expressed opinion second.
@@
This robustness with respect to $\alpha$ is lost in the model without self-anticonformity (bottom row in Fig. 3). In this variant, the stationary curves $c_{S}(p)$, $c_{\sigma}(p)$, and $d(p)$ vary with $\alpha$. An intuitive explanation follows from the fact that private conformity is conditional on the expressed state in this model: even if the $q$-panel is unanimous, a private update is allowed only when the unanimous opinion of the panel coincides with the target’s expressed opinion $S_{i}$. Hence, the expressed layer effectively acts as a filter that can block private conformity events. As a consequence, changing $\alpha$ affects not only the relative frequency of updates at the private and expressed levels, but also the effectiveness of private conformity. Varying $\alpha$ can modify not only the location of the critical point $p^{*}$ but also the type of phase transition. For $q\leq 2$ the transition remains continuous, and for $q\geq 4$ it remains discontinuous, as in Kamińska and Sznajd-Weron (2025). However, for $q=3$ we observe a qualitative change: there exists a critical value $0.8\leq\alpha^{*}\leq 0.9$ above which the discontinuous transition becomes continuous.

### v3

# Competition between private and expressed opinions in binary choice: the $\alpha$-EPO $q$-voter model

@@ Abstract
People often express opinions that differ from their privately held views, a phenomenon known in economy as preference falsification. Expressed-private opinion (EPO) models capture this by assigning each agent two dynamical variables: a private (internal) and an expressed (external) opinion. Within the nonlinear $q$-voter model, two EPO variants have been studied so far: with and without self-anticonformity. In both formulations, agents update private and expressed binary opinions, one after another and at the same rate, which has led to two update schemes studied previously: AT (act then think), in which an agent first updates its expressed and then its private opinion, and TA (think then act), in which the order is reversed. To eliminate this ad hoc distinction and quantify the interplay between private and expressed opinions, we introduce the $\alpha$-EPO $q$-voter model with asynchronous updating – in each elementary step, an agent updates its private opinion with probability $\alpha$ or its expressed opinion with complementary probability $1-\alpha$. We derive mean-field theory and, for the first time for EPO $q$-voter dynamics, a pair approximation, and validate them with Monte Carlo simulations on artificial and real organizational networks. Comparing the two model variants, we show that the collective outcome controlled by $\alpha$ strongly depends on self-anticonformity: with self-anticonformity the results are robust to $\alpha$, whereas without it $\alpha$ shifts the agreement–disagreement threshold and, for $q=3$, changes the phase transition from continuous to discontinuous. Based on these results, we briefly discuss implications for real societies that may be relevant to managers and policymakers.
@@ I Introduction
It is well known that people do not always express their true preferences, a phenomenon known as preference falsification, because they may fear social disapproval, sanctions, or negative consequences arising from openly stating views that deviate from perceived social norms [20]. A natural modeling approach to such situations is to distinguish between an agent’s private and expressed states. This idea underlies the class of expressed-private opinion (EPO) models [4]; for a short and recent review, see [16].
@@
Here, we focus on a specific aspect of these models: the effect of unequal update rates for private and expressed opinions in a binary choice framework, which is relevant to surprisingly complex problems [35]. We study this effect within the nonlinear $q$-voter model, one of the most popular opinion-dynamics models introduced in physics [3, 30, 28, 11]. There are various approaches to modeling the interplay between private and expressed (often called public) opinions. Some models assume that private opinions do not change over time [26, 8, 25, 24]. By contrast, others allow both behavior and attitude to change at the same rate [1, 37, 21, 12, 4, 31]. Finally, there are models in which the update rates for expressed and private opinions differ. An interesting example is the model proposed by Piras et al. [33], in which the agents are generally more likely to change their behavior than their attitudes. However, when agents experience strong cognitive dissonance [6], whether they change their behavior or opinion depends on how often they have altered their opinions in the past. Another example, much closer to what we propose in this paper, is the concealed voter model [9], in which agents change their expressed opinion by copying the opinion of a randomly selected neighbor (with probability $c$) or externalizing their internal opinion (with probability $e$), while the private opinion is updated by internalizing the expressed one (with probability $i$).
@@
Despite the fact that $q$-voter models [3, 30, 28, 11] typically follow an asynchronous updating scheme – at each elementary time step a single voter updates its opinion – the two $q$-voter-based EPO models [13, 16] assume that private and expressed opinions are updated in a quasi-synchronous manner. In this formulation, at each step a single voter updates both its private and public opinion in a specified order. Within this updating scheme, it has been shown that the stationary average private and expressed opinions (i.e., the collective-level aggregates) are independent of whether agents update their private opinion before their expressed opinion, or vice versa [13, 16]. It remains unclear, however, whether introducing rate asymmetry between private and expressed opinions alters the collective dynamics. We therefore extend these models by introducing an asynchronous update scheme, in which only one layer is updated at each step: private opinions are updated with probability $\alpha$, and expressed opinions with the complementary probability $1-\alpha$. We refer to these extensions as $\alpha$-EPO models. We study both variants considered in  [13, 16] and analyze the resulting dynamics on synthetic and empirical networks using mean-field and pair approximation, as well as Monte Carlo simulations.
@@
Studying both variants in parallel is particularly interesting because, as recently suggested, the older version of the model can be interpreted as incorporating a form of self-anticonformity [13], whereas the newer version does not include this mechanism [16]. Furthermore, Jędrzejewski et al. [13] and Kamińska and Sznajd-Weron [16] considered their models only on a complete graph, corresponding to a fully connected society in which everyone knows everyone. For large social systems, this assumption is clearly unrealistic. Here, we move beyond the complete graph by studying $\alpha$-EPO $q$-voter models on networks using Monte Carlo (MC) simulations and pair approximation (PA) analysis.
@@ II The model
We consider a system of $N$ agents, indexed by $i=1,\ldots,N$, each described by two binary variables: $S_{i}(t)\in\{-1,+1\}$, the expressed opinion of agent $i$ at time $t$ (visible to other agents), and $\sigma_{i}(t)\in\{-1,+1\}$, the private opinion of agent $i$ at time $t$ (not visible to others). The binary variables represent a choice between two alternatives, such as yes or no; for or against; adopt or not adopt, etc. Agents occupy the nodes of an undirected, unweighted network, whose topology specifies the interaction patterns within the system. For each agent $i$, let $\mathcal{N}(i)$ denote the set of its neighbors and let $k_{i}=|\mathcal{N}(i)|$ be the degree of node $i$; thus, agent $i$ can interact only with agents in $\mathcal{N}(i)$. In conformity events, the influence group (the $q$-panel) is formed by selecting uniformly at random, without replacement, $q$ neighbors of agent $i$, i.e.,a subset $\mathcal{Q}\subset\mathcal{N}(i)\;\wedge\;|\mathcal{Q}|=q$. In all analytical calculations and algorithmic description of the model we assume that every node in the network has at least $q$ neighbors.
@@
Agents can update their opinions at both levels (expressed and private) due either to social influence from their neighbors or to the need to reduce internal dissonance, which arises when private and expressed opinions are in conflict. The updating rules for public and private opinions differ; however, at both levels we consider two mechanisms: independence and conformity. Independence at the expressed level means that an agent simply verbalizes its private opinion, i.e., its expressed opinion is set equal to the private one, thereby reducing the agent’s internal dissonance. In contrast, independence at the private level corresponds to stochastic noise, as in many other versions of the nonlinear voter model [2, 17, 18, 32, 36]. Specifically, we use the formulation proposed in [30]: under independence, an agent flips its private opinion to the opposite one with probability $1/2$.
@@
Conformity at both levels is driven by social pressure exerted by other agents. In our models, the source of this social influence is a group of $q$ randomly selected neighbors, called the $q$-panel. At the expressed level, conformity comprises two scenarios, depending on the agent’s initial state. If an agent is initially in harmony (i.e., its private and expressed opinions coincide), it is less susceptible to influence. Therefore, it adjusts its expressed opinion to match the group only if the $q$-panel is unanimous, that is, if all $q$ selected neighbors share the same expressed opinion. This type of behavior is called compliance [29]. By contrast, an agent who is in dissonance and thus experiences internal tension is more willing to change. In this case, it is enough that at least one member of the $q$-panel publicly expresses an opinion consistent with the agent’s private beliefs to encourage the agent to speak up, align its expressed opinion with its private one, and thereby relieve dissonance. This type of response to social influence is referred to as disinhibitory contagion [29], a phenomenon in which breaking the norm by even a single individual incites others to act according to their true intentions.
@@
Conformity at the private level also requires a source of influence. Here, a key difference emerges between the models proposed by Jędrzejewski et al. [13] and Kamińska and Sznajd-Weron [16], and this distinction carries over to the two variants considered in this paper. In the first case, following Jędrzejewski et al. [13], an agent conforms to a unanimous $q$-panel regardless of its expressed opinion. Consequently, the agent may change its private opinion to match the group while still expressing the opposite view. We refer to this as self-anticonformity, since the agent, at the expressed level, disagrees not only with the group but also with its own private opinion. In the second variant, following Kamińska and Sznajd-Weron [16], this scenario is prohibited. In other words, an agent conforms privately only if the $q$-panel is unanimous and its opinion matches the agent’s expressed one. This also implies that conformity on the private level cannot lead to dissonance.
@@
In this paper, we introduce an extension that applies to both variants of the model described above, allowing for a more flexible relationship between updates at the expressed and private levels. The proposed extension captures the fact that a change in private opinion is not necessarily followed by a change in behavior, and vice versa: behavior may change without altering beliefs [19]. Both Jędrzejewski et al. [13], Kamińska and Sznajd-Weron [16] assumed that during a single update, both expressed and private opinions are updated. They considered two update schemes: “act then think” (AT), in which the expressed opinion is updated first and the private one second, and “think then act” (TA), in which the private opinion is updated first and the expressed one second. Here, we assume that with probability $\alpha$ an agent updates its private opinion, and with complementary probability $1-\alpha$ it updates its expressed opinion. The algorithm for an elementary update of the $\alpha$-EPO model, taking time $\Delta t$, is described by Algorithm 1 and illustrated in Figs. 1 and 2.
@@
8:   Draw random set $\mathcal{Q}\subset\mathcal{N}(i)$ such that $|\mathcal{Q}|=q$
@@
28:   Draw random set $\mathcal{Q}\subset\mathcal{N}(i)$ such that $|\mathcal{Q}|=q$
@@
Let us, following [13, 16], use $\uparrow$ to denote the opinion +1+1 and $\downarrow$ to denote the opinion 1-1. To describe the state of a single agent at time $t$, we use the ordered pair $(S(t),\sigma(t))$, where the first element is the expressed opinion, followed by the private one. Accordingly, the set of possible agent states is $X=(S,\sigma)\in\{\uparrow\uparrow,\uparrow\downarrow,\downarrow\uparrow,\downarrow\downarrow\}$. The macroscopic state of the system at time $t$ can be described by the fractions of agents in each state, denoted by $c_{X}(t)$, defined as
@@
To enable comparison with Jędrzejewski et al. [13] and Kamińska and Sznajd-Weron [16] instead of analyzing the concentrations of agents in all possible states separately, we consider the following quantities:
@@ III.1 Monte Carlo simulations
We use Monte Carlo simulations both to validate our analytical calculations for complete and random graphs, as well as to extend the analysis to other topologies, including the empirical networks. To obtain smooth trajectories comparable with the mean-field and pair approximations (which effectively correspond to the thermodynamic limit), we simulate relatively large systems with $N=10^{4}$ and $N=10^{5}$ nodes. We also conduct MC simulations on empirical networks from organizations [7]. In the latter case, some nodes may have degree lower than $q$. In such cases we assume that the $q$-panel consists of all neighbors, i.e., for each agent $i$ we take $q_{i}=\min(k_{i},q)$.
@@ III.2.1 Model with self-anticonformity
We can proceed analogously to Jędrzejewski et al. [13], that is we add (14) and (15), then (14) and (16). Then, using Eqs. (3) and (4) we finally obtain
@@
Hence for $\alpha\neq 0$ and $\alpha\neq 1$ we obtain exactly the same system of equations as Jędrzejewski et al. [13], namely
@@ III.3 Pair approximation
Pair approximation is an analytical approach that relaxes the assumption of a completely homogeneous agent system by accounting for dynamical correlations between pairs of agents in different states. It has been successfully applied to many versions of the $q$-voter model, in which agents are described by a single opinion [15, 11, 36, 14, 34], but not to the EPO $q$-voter models. Within this framework, the evolution equations for the state concentrations in the model with self-anticonformity are given by the following system (for the model without self-anticonformity, the framed terms are omitted):
@@
As in other homogeneous pair-approximation treatments of $q$-voter models [14, 23], correlations between node degree and state do not emerge unless they are present in the initial condition. Hence, we can set $\langle k\rangle_{X}=\langle k\rangle$ for all $X$ and extract the opinion concentration from Eq. (42) according to:
@@ IV.1 Mean-field results
As shown in Fig. 3, both models exhibit a phase transition between an ordered phase (agreement), in which the majority of agents share the same opinion ($c_{S}\neq 0.5$, $c_{\sigma}\neq 0.5$), and a disordered phase (disagreement), in which opinions are evenly split ($c_{S}=0.5$, $c_{\sigma}=0.5$). The location of the critical point $p^{*}$ and the type of the transition differ between the two $\alpha$-EPO variants. Nevertheless, within each variant, $c_{S}(p)$ and $c_{\sigma}(p)$ undergo the transition at the same critical point, and the transition is of the same type for both quantities. In the ordered phase ($p<p^{*}$), the majority opinion is more pronounced at the expressed level than at the private level. This result is consistent with pluralistic ignorance, i.e., a mismatch in which privately held views are more diverse than publicly expressed ones because individuals misperceive the prevailing norm and adjust their public behavior accordingly [27].
@@
The role of $\alpha$ differs qualitatively between the two variants of the $\alpha$-EPO model. In the version with self-anticonformity, the stationary phase diagrams $c_{S}(p)$ and $c_{\sigma}(p)$ are independent of $\alpha$ and therefore coincide with those reported in Jędrzejewski et al. [13] (top row in Fig. 3). This is consistent with Eqs. (20)–(21), which, for $0<\alpha<1$, do not contain $\alpha$. In this variant, $\alpha$ affects only the stationary dissonance $d$, which increases with $\alpha$. This agrees with Jędrzejewski et al. [13], where dissonance was also the only quantity that distinguished the two update schemes: act then think (AT), in which the expressed opinion was updated first and the private opinion second, and think then act (TA), in which the private opinion was updated first and the expressed opinion second.
@@
This robustness with respect to $\alpha$ is lost in the model without self-anticonformity (bottom row in Fig. 3). In this variant, the stationary curves $c_{S}(p)$, $c_{\sigma}(p)$, and $d(p)$ vary with $\alpha$. An intuitive explanation follows from the fact that private conformity is conditional on the expressed state in this model: even if the $q$-panel is unanimous, a private update is allowed only when the unanimous opinion of the panel coincides with the target’s expressed opinion $S_{i}$. Hence, the expressed layer effectively acts as a filter that can block private conformity events. As a consequence, changing $\alpha$ affects not only the relative frequency of updates at the private and expressed levels, but also the effectiveness of private conformity. Varying $\alpha$ can modify not only the location of the critical point $p^{*}$ but also the type of phase transition. For $q\leq 2$ the transition remains continuous, and for $q\geq 4$ it remains discontinuous, as in Kamińska and Sznajd-Weron [16]. However, for $q=3$ we observe a qualitative change: there exists a critical value $0.8\leq\alpha^{*}\leq 0.9$ above which the discontinuous transition becomes continuous.